# Enumerating Finite Braid Group Orbits on $SL_{2}(\mathbb{C})$-Character Varieties

### v2

@@ 1.1 Prior Work
MCG-finite representations correspond to algebraic solutions to nonlinear differential equatons, such as the Painlevé VI equation and more generally the Schlesinger equations [7]. Regarded as a nonlinear analogue of the hypergeometric equation, the Painlevé VI (PVI) equation posed over 100 years ago, has inspired much work on the study of finite orbits. Algebraic solutions to the PVI equation correspond to irreducible MCG-finite representations when $n=2$ and $T+1=4$ and motivates progress towards a nonlinear Schwartz list [3]. Among the many advances made in this area, notable contributions on finite orbits come from Hitchin [17], Dubrovin-Mazzocco [12], Boalch [2], Kitaev [19] and an eventual computer-aided classification by Lisovy-Tykhyy [22]. The history and further details regarding the 45 equivalence classes of exceptional algebraic solutions of PVI can be found in [1]. Research on finite orbits of the character variety is still ongoing, see [25], [5], and more recently [4].
@@
The work of Dubrovin-Mazzocco originally studied the PVI equation in connection with reflection groups [12]. Expanding on this idea, solutions constructed from finite reflection groups through special cases of middle convolution, like the “Okamoto transform,” were explored by others in [2], [23], [10]. More generally, a computer-aided classification was provided by Tykhyy in [31], with explicit orbits listed when $T+1=4,5$ and conjectural descriptions of the different possibilities for $T+1\geq 6$. We found exact matches for orbits given in this list, except in two cases where the closest match differs in just one choice of parameter. As both these examples almost match with the same orbit they provided, we suspect this difference might be due to a typo. Details on these two tuples are provided in Section 5.1.3 and Section 5.1.4 and the differing coordinates are marked with an asterisk in the tables.
@@
The connection between finite complex reflection groups and middle convolutions was further explored more recently in the work of Lam-Landesman-Litt [20, Cor. 1.1.7] who proved that all MCG-finite tuples with at least one matrix of infinite order are either pullback representations (classified in [11]) or come from finite complex reflection groups via middle convolution.
@@
Our work investigates the finite complex reflection groups for MCG-finite tuples and examines their middle convolution in relation to the braid group orbit. In the applicable cases, we also explore how the examples we find relate to the classifications provided by Lisovy-Tykhyy [22] and Tykhyy [31]. Note that the analysis in these two papers exclude the tuples where all the matrices share a common eigenvector; these correspond to non-irreducible cases and were classified by Cousin-Moussard [14, Section 3].
@@ 2.1 Middle Convolution
Defined by Katz [18], middle convolution is an invertible functor in the category of perverse sheaves depending on some parameter $\lambda\in\mathbb{C}^{\times}$ and is denoted $MC_{\lambda}$. Middle convolution can also be viewed as a functor in the category of local systems on the punctured sphere to itself, and with this interpretation, Katz proved that all irreducible rigid local systems on the punctured sphere can be constructed from rank 1 local systems via scaling or middle convolution.
Letting $k$ be a field and $F_{T}$ denote the free group on $T$ generators, middle convolution can equivalently be understood as a functor on the category of finite dimensional $k[F_{T}]$ modules to itself as described by Dettweiler-Reiter [8, Prop. 2.6]. The relation to local systems on the $T+1$-punctured sphere is by taking $k=\mathbb{C}$ and viewing a tuple of $T$ matrices as defining the action of $\mathbb{C}[F_{T}]$ on the module i.e. a representation of $\mathbb{C}[F_{T}]$. In this way irreducible tuples correspond to irreducible representations. While middle convolution preserves irreducibility [8, Cor. 3.6], it typically changes the dimension of the module (rank of the local system on $T+1$-punctured sphere) and does not preserve basic properties e.g. finiteness of the subgroup generated by the matrices (i.e. monodromy group). Important to the study of finite orbits, one desirable property of middle convolution is that it is equivariant with the action of the braid group $B_{T}$ [8, Theorem 5.1]. One way to construct nontrivial finite orbits is to take irreducible tuples generating a finite subgroup, hence finite braid group orbit, and taking the middle convolution to (hopefully) get a tuple generating an infinite group which will necessarily be irreducible and have finite braid group orbit.
The purely algebraic analogue by Dettweiler-Reiter [8] gives an explicit description of middle convolution in terms of tuples of matrices and relates some properties of middle convolution to the matrices in the tuple.
@@
Given a tuple of $T$ matrices $\textbf{A}=[A_{1},A_{2},\dots,A_{T}]$ with $A_{i}\in GL_{n}(\mathbb{C})$, the $\mathbf{product}$ of such a tuple is the product of the matrices in the tuple, in the ordering given by the tuple. So the product of the tuple $\mathbf{A}$ is $A_{1}A_{2}\cdots A_{T}$. The $\mathbf{inverse}$ of the tuple A is the tuple $[A_{T}^{-1},\dots,A_{1}^{-1}]$, denoted $\mathbf{A^{-1}}$. The $\mathbf{inverse}$ $\mathbf{product}$ of a tuple is the matrix product of the inverse tuple.
@@
Let $V$ denote the $\mathbb{C}[F_{T}]$-module defined by the tuple $\textbf{A}=[A_{1},A_{2},\dots,A_{T}]$ and let $\lambda\in\mathbb{C}^{\times}$.
@@
Proofs of this statement, as well as further details and properties of the middle convolution can be found in [9, Section 2] and [8, Section 2-5]. ∎
[…]
@@
Proved in [8, Theorem 5.5]. ∎
Our code for middle convolution uses the formulation given by Dettweiler-Reiter in [9, Definition 2.1] as opposed to their formulation in [8, Section 2] which only differs by changing the order of blocks in the matrices. The code requires a matrix group $W$ containing the matrices in the tuple $\mathbf{A}$ and the base ring of $W$ should contain the parameter $\lambda$ used for middle convolution. Note that the final function $Compute\_MC$ will automatically check if our intermediate constructions act invariantly on the subspaces $K$ and $L$, as defined in the notation of [9, Definition of 2.1]. The full computer code can be found online at amalvayalin.github.io/research.html.
@@ 2.2 Finite Complex Reflection Groups
A real reflection can be identified as a matrix with one eigenvalue -1 and all other eigenvalues are 1, and pseudo-reflections refers to matrices with one eigenvalue a root of unity and all other eigenvalues 1. A finite complex reflection group (FCRG) is a finite group generated by pseudo-reflections. As we are only working with complex groups, the distinction between real or complex eigenvalues is not relevant and henceforth reflections can refer to real reflections or pseudo-reflections.
All irreducible finite complex reflection groups were classified by Shephard-Todd [26], as belonging to an infinite family of imprimitive reflection groups or being one of 34 exceptional primitive cases.
@@
The rank of a reflection group is the smallest dimension of the vector space on which the group acts irreducibly. Every finite complex reflection group of rank $n$ has a minimal generating set consisting of either $n$ or $n+1$ reflections. The group is said to be well-generated if it can be generated by $n$ reflections.
As we want irreducible MCG-finite T+1-tuples, we know that our nice $T$-tuple (see Defintion 3.0.1) must at least be of length $n$ to generate an irreducible FCRG, so in each group we only need to search for $T$-tuples with $T>n$. Among the rank three and four primitive irreducible groups, only $G_{31}$ is not well-generated (see Cartan matrices in [21, Appendix Section 3]. An imprimitive group $G(m,p,n)$ is well-generated if $p=m$ or $p=1$.
Associated to each FCRG of rank $n$ is a set of $n$ invariants called the degrees of the reflection group [26, Theorem 5.1], [21, Table D.3]. An important property of these invariants relevant for our purposes is their relation to the eigenvalues of the group elements. Note that we already know the eigenvalues for any element of the group must be roots of unity since the group is finite.
@@ 2.3 Representations and Character Varieties
In the case of the punctured sphere $\Sigma_{0,T+1}$, the fundamental group is $\pi_{1}(\Sigma_{0,T+1})=\langle\gamma_{1},\dots,\gamma_{T+1}\mid\gamma_{1}*\gamma_{2}*\cdots*\gamma_{T+1}=1\rangle$. Any representation of $\Gamma=\pi_{1}(\Sigma_{0,T+1})$ is identified by $\rho(\gamma_{i})=M_{i}\in G$ such that $M_{1}\cdot M_{2}\cdots M_{T+1}=I$ is the identity matrix. Two such representations $\rho$ and $\rho^{\prime}$ are equivalent if the tuple $[\rho(\gamma_{1}),\dots,\rho(\gamma_{T+1})]$ is related to $[\rho^{\prime}(\gamma_{1}),\dots,\rho^{\prime}(\gamma_{T+1})]$ by simultaneous conjugation.
Taking $G=SL_{2}(\mathbb{C})$, we can consider the $SL_{2}(\mathbb{C})$-character variety of $\pi_{1}(\Sigma_{0,T+1})$.
Given a tuple of of matrices $[M_{1},M_{2},\dots,M_{T+1}]$, let
@@
Let $[M_{1},M_{2},\dots,M_{T+1}]\in SL_{2}(\mathbb{C})^{T+1}$ such that $\Pi M_{i}=I$. Then up to simultaneous conjugation, any such $(T+1)$-tuple can be identified by the following trace coordinates:
@@
For proof and further discussion of character varieties of the punctured sphere see [15, Section 4] and [24, Section 5]. ∎
Such trace coordinates are also sometimes referred to as the $\mathbf{signature}$ of the tuple.
@@ 3 Methods
Since the product of $[A_{1},\cdots,A_{T+1}]$ must be the identity, we know any MCG-finite tuple is determined by the first $T$ matrices, up to simultaneous conjugation. As middle convolution does not preserve products (see Remark 2.1.3), our idea is to take middle convolution of the first $T$ matrices to get $2\times 2$ matrices and then induce a representation of $\pi_{1}(\Sigma_{0,T+1})$ by taking the $T+1$-th matrix to be the inverse product. The condition for middle convolution being rank 2, (stated in Theorem 2.1.4) depends on the dimension of the subspace fixed by, or equivalently the multiplicity of the eigenvalue 1 of each matrix in the tuple and the eigenvalues of their product. For much better control on the former, we consider reflections since they only have one nontrivial eigenvalue, of multiplicity 1.
Thus, for each irreducible finite complex reflection group $W$ of rank $n$, we want to find all irreducible $T$-tuples of reflections $\textbf{A}=[A_{1},A_{2},\dots,A_{T}]$ in $W$ such that for a suitable choice of parameter $\lambda$, $MC_{\lambda}(\textbf{A})$ has rank 2. Given $T$, the formula given in Theorem 2.1.4 tells us that the product $\Pi_{i=1}^{T}A_{i}$ of reflections in A should have a nontrivial eigenvalue $\lambda^{-1}$ of multiplicity $T-2$ so that $MC_{\lambda}(\textbf{A})=[\tilde{A_{1}},\tilde{A_{2}},\dots,\tilde{A_{T}}]$ is a $T$-tuple of rank 2. We give a name to tuples satisfying these conditions.
Definition 3.0.1.
A $T$-tuple $\textbf{A}=[A_{1},A_{2},\dots,A_{T}]$ of matrices from a finite complex reflection group $W$ is called $\mathbf{nice}$ if A satisfies the following:
  1. 1.
    All $A_{i}$ are reflections in $W$
  2. 2.
    $W=\langle A_{1},A_{2},\dots,A_{T}\rangle$
  3. 3.
    $(A_{1}A_{2}\cdots A_{T})^{-1}$ has an eigenvalue $\lambda\neq 1$ of multiplicity $T-2$.
The condition of generating the whole group is to ensure that we are not “over-counting” as some reflection groups contain other reflection groups as subgroups, and we only want to study irreducible MCG-finite tuples.
Remark 3.0.2.
Given a nice $T$-tuple $\mathbf{A}=[A_{1},A_{2},\dots,A_{T}]$, we know there exists some $\lambda\neq 1$ such that $MC_{\lambda}(\mathbf{A})=[\tilde{A_{1}},\tilde{A_{2}},\dots,\tilde{A_{T}}]$ is a $T$-tuple of rank 2. taking the $T+1$-th matrix to be
$\widetilde{A_{T+1}}=(\Pi_{i=1}^{T}\tilde{A_{i}})^{-1}$
we get a $T+1$-tuple $[\tilde{A_{1}},\tilde{A_{2}},\dots,\tilde{A_{T}},\widetilde{A_{T+1}}]$ whose product is the identity. As $\mathbf{A}$ comes from a finite group we know it has finite braid group orbit, and since middle convolution is equivariant with the braid group action (see Section 2.1), we know that $[\tilde{A_{1}},\tilde{A_{2}},\dots,\tilde{A_{T}},\widetilde{A_{T+1}}]$ must be MCG-finite. Moreover, if $\{A_{1},A_{2},\dots,A_{T}\}$ generates an irreducible group, then as middle convolution preserves irreducibility, we know our $T+1$-tuple is irreducible as well. Thus, we have used a nice $T$-tuple to produce an irreducible MCG-finite $T+1$-tuple.
Remark 3.0.3.
For some nice $T$-tuples, it is possible that multiple choices of $\lambda$ have multiplicity $T-2$ and hence the different choices of parameter $\lambda$ gives rise to possibly different rank 2 MCG-finite $T+1$-tuples.
Remark 3.0.4.
The eigenvalue condition of multiplicity $T-2$ tells us that in any complex reflection group of rank $n$, for $T>n+2$ the middle convolution will not be rank 2 so there are no nice $T$-tuples. By Definition 2.2.1, since we need at least $n$ reflections to generate the whole group, we know there are no nice tuples when $T<n$. For $T=n+2$, the product of the $T$ matrices would be a scalar matrix (as there must be an eigenvalue of multiplicity $n$) which we can take to be the identity, reducing to the $T=n+1$ case. So for each $n\in\{3,4\}$, we only search the primitive reflection groups for nice tuples of length $T=n$ or $T=n+1$.

### v3

@@ 1.2 Prior Work
MCG-finite representations correspond to algebraic solutions to nonlinear differential equations, such as the Painlevé VI equation and more generally the Schlesinger equations [cou17]. Regarded as a nonlinear analogue of the hypergeometric equation, the Painlevé VI (PVI) equation, posed over 100 years ago, has inspired much work on the study of finite orbits. Algebraic solutions to the PVI equation correspond to irreducible MCG-finite representations when $n=2$ and $T+1=4$ and motivates progress towards a nonlinear Schwartz list [boalch2007towards]. Among the many advances made in this area, notable contributions on finite orbits come from Hitchin [hitchinponcelet], Dubrovin-Mazzocco [dubromazz2000], Boalch [boalch2005klein], Kitaev, [kitaev2006] and an eventual computer-aided classification by Lisovyy-Tykhyy [T-L]. The history and further details regarding the 45 equivalence classes of exceptional algebraic solutions of PVI can be found in [P6survey]. Research on finite orbits of the character variety is still ongoing, see [Gol2024], [calli2018], and more recently [bronsteinmaret].
@@
The work of Dubrovin-Mazzocco originally studied the PVI equation in connection with reflection groups [dubromazz2000]. Expanding on this idea, solutions constructed from finite reflection groups through special cases of middle convolution, like the “Okamoto transform,” were explored by others in [boalch2005klein], [okamototrans], [dettweilerpainleve]. More generally, a computer-aided classification was provided by Tykhyy in [Tyk], with explicit orbits listed when $T+1=4,5$ and conjectural descriptions of the different possibilities provided for $T+1\geq 6$. We found exact matches for orbits given in this list, except in two cases where the closest match differs in just one choice of parameter. As that parameter is considered up to sign, we believe our orbits still belong to those close-match orbits. Details on these two tuples are provided in Section 5.1.3 and Section 5.1.4, and the differing coordinates are marked with an asterisk in the tables.
@@
The connection between finite complex reflection groups and middle convolutions was further explored more recently in the work of Lam-Landesman-Litt [LLL, Cor. 1.1.7], who proved that all MCG-finite tuples with at least one matrix of infinite order are either pullback representations (classified in [diarra2013]) or come from finite complex reflection groups via middle convolution.
@@
Our work investigates the finite complex reflection groups for MCG-finite tuples and examines their middle convolution in relation to the braid group orbit. In the applicable cases, we also explore how the examples we find relate to the classifications provided by Lisovyy-Tykhyy [T-L] and Tykhyy [Tyk]. Note that the analysis in these two papers excludes the tuples where all the matrices share a common eigenvector; these correspond to non-irreducible cases and were classified by Cousin-Moussard [cousinmouss, Section 3].
Released shortly after our own work was posted to arXiv, theoretical results by Bronstein-Maret [bronsteinmaret] completed the classification of finite orbits by studying the case where all local monodromies have finite order. In light of their research and the work before them, the tuples we provide serve as concrete examples of representations with finite orbits.
@@ 2.1 Middle Convolution
Defined by Katz [katz1996], middle convolution is an invertible functor in the category of perverse sheaves depending on some parameter $\lambda\in\mathbb{C}^{\times}$ and is denoted $MC_{\lambda}$. Middle convolution can also be viewed as a functor in the category of local systems on the punctured sphere to itself, and with this interpretation, Katz proved that all irreducible rigid local systems on the punctured sphere can be constructed from rank 1 local systems via iteratively applying scaling and middle convolution operations.
Letting $k$ be a field and $F_{T}$ denote the free group on $T$ generators, middle convolution can equivalently be understood as a functor on the category of finite dimensional $k[F_{T}]$ modules to itself as described by Dettweiler-Reiter [KatzAlgo, Prop. 2.6]. The relation to local systems on the $T+1$-punctured sphere is by taking $k=\mathbb{C}$ and viewing a tuple of $T$ matrices as defining the action of $\mathbb{C}[F_{T}]$ on the module i.e. a representation of $\mathbb{C}[F_{T}]$. In this way irreducible tuples correspond to irreducible representations, which correspond to local systems. While middle convolution preserves irreducibility [KatzAlgo, Cor. 3.6], it typically changes the dimension of the module (rank of the local system on the $T+1$-punctured sphere) and does not preserve basic properties e.g. finiteness of the subgroup generated by the matrices, or equivalently, the monodromy group of the associated local system. Important to the study of finite orbits, one desirable property of middle convolution is that it is equivariant with the action of the braid group $B_{T+1}$ [KatzAlgo, Theorem 5.1]. One way to construct nontrivial finite orbits is to take irreducible tuples generating a finite subgroup, hence finite braid group orbit, and taking the middle convolution to (hopefully) get a tuple generating an infinite group which will necessarily be irreducible and have finite braid group orbit.
The purely algebraic analogue by Dettweiler-Reiter [KatzAlgo, Definition 2.5] gives an explicit description of middle convolution in terms of tuples of matrices and relates some properties of middle convolution to the matrices in the tuple.
@@
Given a tuple of $T$ matrices $\textbf{A}=[A_{1},A_{2},\dots,A_{T}]$ with $A_{i}\in GL_{n}(\mathbb{C})$, the $\mathbf{product}$ of such a tuple is the product of the matrices in the tuple, in the ordering given by the tuple. So the product of the tuple $\mathbf{A}$ is $A_{1}A_{2}\cdots A_{T}$. The $\mathbf{inverse}$ of the tuple A is the tuple $A_{T}^{-1},\dots,A_{1}^{-1}]$, denoted $\mathbf{A^{-1}}$. The $\mathbf{inverse}$ $\mathbf{product}$ of a tuple is the matrix product of the inverse tuple.
@@
Let $V$ denote the $\mathbb{C}[F_{T}]$-module defined by the irreducible tuple $\textbf{A}=[A_{1},A_{2},\dots,A_{T}]$ and let $\lambda\in\mathbb{C}^{\times}$.
@@
    $\dim(MC_{\lambda}(V))=\sum_{i=1}^{T}\operatorname{rk}(A_{i}-I)-(\dim(V)-\operatorname{rk}(\lambda\cdot A_{1}A_{2}\cdots A_{T}-I))$
Proof.
Proof of this statement, as well as further details and properties of the middle convolution can be found in [MC-L, Section 2] and [KatzAlgo, Section 2-5]. ∎
Remark 2.1.5.
@@
In light of Theorem 2.1.4, to get an MCG-finite $T+1$-tuple of rank 2 we want to find a $T$-tuple of reflections (hence $\operatorname{rk}(A_{i}-I)=1$) such that the tuple product has a nontrivial eigenvalue of multiplicity T2T-2. This is why we search for nice tuples, see Definition 1.1.2.
Theorem 2.1.6.
(Dettweiler-Reiter, 2007) Let $\lambda\in\mathbb{C}^{\times}$, and $MC_{\lambda}(A_{1},\dots,A_{T})=[\tilde{A_{1}},\dots,\tilde{A_{T}}$]. Let $A_{T+1}$, $\tilde{A}_{T+1}$ be the inverse product of the tuples $\mathbf{A}$ and $\mathbf{\tilde{A}}$ respectively. Writing $J(\alpha,l)$ for the Jordan block of size $l$ with eigenvalue $\alpha$, the eigenvalues before and after middle convolution are related as follows:
  1. (i)
    Every Jordan block $J(\alpha,l)$ in the Jordan decomposition of $A_{k}$ contributes a Jordan block $J(\alpha\lambda,l^{\prime})$ to the Jordan decomposition of $\tilde{A_{k}}$ where:
    $l^{\prime}:=\begin{cases}l&\text{if $\alpha\neq 1,\lambda^{-1}$}\\ l-1&\text{if $\alpha=1$}\\ l+1&\text{if $\alpha=\lambda^{-1}$}\end{cases}$
    The only other Jordan blocks in the Jordan decomposition of $\tilde{A_{k}}$ is $J(1,1)$.
  2. (ii)
    Every Jordan block $J(\alpha,l)$ in the Jordan decomposition of $A_{T+1}=(A_{1}A_{2}\cdots A_{T})^{-1}$ contributes a Jordan block $J(\alpha\lambda^{-1},l^{\prime})$ to the Jordan decomposition of $\tilde{A}_{T+1}=(\tilde{A_{1}}\tilde{A_{2}}\cdots\tilde{A_{T}})^{-1}$ where:
    $l^{\prime}:=\begin{cases}l&\text{if $\alpha\neq 1,\lambda$}\\ l+1&\text{if $\alpha=1$}\\ l-1&\text{if $\alpha=\lambda$}\end{cases}$
    The only other Jordan blocks in the Jordan decomposition of $\tilde{A}_{T+1}$ is $J(\lambda^{-1},1)$.
@@
This statement in terms of the Jordan decomposition is a reformulation of [dettweilerpainleve, Lemma 5.1] and follows from [KatzAlgo, Lemma 4.1]. See also [LLL][Prop. 3.2.2]. ∎
@@
Proved in [KatzAlgo, Theorem 5.5]. ∎
Our code for middle convolution uses the formulation given by Dettweiler-Reiter in [MC-L, Definition 2.1] rather than their formulation in [KatzAlgo, Section 2] which only differs by changing the order of blocks in the matrices. The code requires a matrix group $W$ containing the matrices in the tuple $\mathbf{A}$ and the base ring of $W$ should contain the parameter $\lambda$ used for middle convolution. Note that the final function Compute_MCCompute\_MC will automatically check if our intermediate constructions act invariantly on the subspaces $\mathcal{K}$ and $\mathcal{L}$, as defined in the notation of [MC-L, Definition of 2.1]. The full computer code can be found online at amalvayalin.github.io/research.html.
@@ 2.2 Finite Complex Reflection Groups
A real reflection can be identified as a matrix with eigenvalue -1 of multiplicity one and all other eigenvalues equal to 1. Pseudo-reflections refer to matrices with a root of unity as an eigenvalue of multiplicity one and all other eigenvalues equal to 1. A finite complex reflection group (FCRG) is a finite group generated by pseudo-reflections. As we are only working with complex groups, the distinction between real or complex eigenvalues is not relevant and henceforth reflections can refer to real reflections or pseudo-reflections.
All irreducible finite complex reflection groups were classified by Shephard-Todd [stfinite], as belonging to an infinite family of imprimitive reflection groups or being one of 34 exceptional primitive cases. These groups come with a canonical finite-dimensional representation in terms of an action generated by reflections on a vector space $V$. Every FCRG of rank $n$, defined below, has a minimal generating set consisting of either $n$ or $n+1$ reflections.
@@
The rank of a reflection group, presented as being generated by reflections on a vector space $V$, is the dimension of $V$. A rank $n$ group is said to be well-generated if it can be generated by $n$ reflections.
As we want irreducible MCG-finite TT+1-tuples, we know that our nice $T$-tuple (see Defintion 1.1.2) must at least be of length $n$ to generate an irreducible FCRG, so in each group we only need to search for $T$-tuples with $T>n$. Among the rank three and four primitive irreducible groups, only $G_{31}$ is not well-generated (see Cartan matrices in [ld2009, Appendix Section 3]. An imprimitive group $G(m,p,n)$ is well-generated if $p=m$ or $p=1$.
Associated to each FCRG of rank $n$ is a set of $n$ invariants called the degrees of the reflection group [stfinite, Theorem 5.1], [ld2009, Table D.3]. An important property of these invariants relevant for our purposes is their relation to the eigenvalues of the group elements. Note that we already know the eigenvalues for any element of the group must be roots of unity since the group is finite.
@@ 2.3 Representations and Character Varieties
In the case of the punctured sphere $\Sigma_{T+1}$, the fundamental group is
$\pi_{1}(\Sigma_{T+1},x)=\langle\gamma_{1},\dots,\gamma_{T+1}\mid\gamma_{1}\cdot\gamma_{2}\dots\gamma_{T+1}=1\rangle$
Any representation of $\Gamma=\pi_{1}(\Sigma_{T+1})$ is identified by $\rho(\gamma_{i})=M_{i}\in G$ such that $M_{1}\cdot M_{2}\cdots M_{T+1}=I$ is the identity matrix. Two such representations $\rho$ and $\rho^{\prime}$ are equivalent if the tuple $\rho(\gamma_{1}),\dots,\rho(\gamma_{T+1})]$ is related to $\rho^{\prime}(\gamma_{1}),\dots,\rho^{\prime}(\gamma_{T+1})]$ by simultaneous conjugation.
Taking $G=SL_{2}(\mathbb{C})$, we can consider the $SL_{2}(\mathbb{C})$-character variety of $\pi_{1}(\Sigma_{T+1})$.
Given a tuple of of matrices $M_{1},M_{2},\dots,M_{T+1}]$, let
@@
Let $M_{1},M_{2},\dots,M_{T+1}]\in SL_{2}(\mathbb{C})^{T+1}$ such that $\Pi M_{i}=I$. Then up to simultaneous conjugation, any such $(T+1)$-tuple can be identified by the following trace coordinates:
@@
For proof and further discussion of character varieties of the punctured sphere see [goldman2009tr, Section 4] and [VM2002, Section 5]. ∎
Definition 2.3.2.
The trace coordinates above are referred to as the $\mathbf{signature}$ of a tuple.
@@ 3 Methods
Since the product of $A_{1},\cdots,A_{T+1}]$ must be the identity, we know any MCG-finite tuple is determined by the first $T$ matrices, up to simultaneous conjugation. As middle convolution does not preserve products (see Remark 2.1.3), our idea is to take middle convolution of the first $T$ matrices to get $2\times 2$ matrices and then induce a representation of $\pi_{1}(\Sigma_{T+1})$ by taking the $T+1$-th matrix to be the inverse product. The condition for middle convolution being rank 2, (stated in Theorem 2.1.4) depends on the dimension of the subspace fixed by, or equivalently the multiplicity of the eigenvalue 1 of each matrix in the tuple and the eigenvalues of their product. For much better control on the former, we consider reflections since they only have one nontrivial eigenvalue, of multiplicity 1.
Thus, for each irreducible finite complex reflection group $W$ of rank $n$, we want to find all irreducible $T$-tuples of reflections $\textbf{A}=[A_{1},A_{2},\dots,A_{T}]$ in $W$ such that for a suitable choice of parameter $\lambda$, $MC_{\lambda}(\textbf{A})$ has rank 2. Given $T$, the formula given in Theorem 2.1.4 tells us that the product $\Pi_{i=1}^{T}A_{i}$ of reflections in A should have a nontrivial eigenvalue $\lambda^{-1}$ of multiplicity T2T-2 so that $MC_{\lambda}(\textbf{A})=[\tilde{A_{1}},\tilde{A_{2}},\dots,\tilde{A_{T}}$] is a $T$-tuple of rank 2. Since $W$ itself is irreducible, if we further required the that $W=\langle A_{1},A_{2},\dots,A_{T}\rangle$, we can avoid “over-counting” as some reflection groups contain other reflection groups as subgroups, and we only want to study irreducible MCG-finite tuples. These conditions are precisely our definition of nice tuples given in Definition 1.1.2!